Kinkead Inc. forecasts that its free cash flow in the coming year, i.e., at t = 1, will be -$10 million, but its FCF at t = 2 will be $20 million. After Year 2, FCF is expected to grow at a constant rate of 4% forever. If the weighted average cost of capital is 14%, what is the firm’s value of operations, in millions?
1
Expert's answer
2014-12-08T10:56:34-0500
At t = 1 FCF = -$10 million, FCF (t = 2) = $20 million. After Year 2 FCF is expected to grow at a constant rate of 4% forever. WACC = 14%. Firm’s value of operations will be: NPV = sum(FCF/(1 + wacc))^n = -10/1.14 + 20/1.14^2 + 20*1.04/(1.14^3*0.14) = 8.58 million.